TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Created by Main.MirkoPojer on 2007-10-09 - 21:28

From Pressure test to Cool-down - Sector 5-6 _ 09/10/2007

JPEG 5-6.jpg:

Feedback of the meeting:

  • Problems with the soldering of the flexible of DFBAJ - To be resolded (after meeting: soldered, waiting for radiography)
  • ElQA on DFBXF.3R5 not yet done (after meeting: possibly done on Thursday)
  • Cool-down could be delayed 1 week (start on w.46)

After meeting:

  • Partial coherence test successfully performed
  • Leak in 15L6 will be repaired on Wednesday; 3 bars will be put during night to check
  • Rail for re-routing of 120 A cables on DFBLE.RR57 is in place; connection on Wednesday

Edit, attach file or add comment

Created by Main.BorisBellesia on 2007-10-09 - 20:16

From Pressure Test to Cool-Down - Tuesday 09 - Point 4


General planning:

  • W41-W42: L5 to 80K
  • W42-W43: Cool-down of the whole sector from 80K up to 4.5K
  • W44: Cool-down of the arc45 from 4.5K up to 1.9K / ElQA of the Circuits of the matching sections / Powering tests of the matching sections
  • W45: ElQA of the circuits of the arc and Powering tests of the matching sections
  • W46: Powering tests of the circuits of the arc

Activitites - w41:

  • Today -09/10 - cooldown L5 to 80K
    • Sector closed since 08/10
    • Access allowed with key and tokens except RE52-UL55
    • See document LHC-S-ES-0017 for particular access conditions
  • ElQA break-down monitoring all along cooldown; AT-MEL
  • QPS activities; AT-MEL
  • CMS vacuum installation/tests; AT-VAC
  • Magnet alignment; TS-SU
  • Cable campaign point5; TS-EL
  • Inner triplet interconnection; AT-MCS

Open Issues + Long term activitites:

  • Swapped cables of RB circuit on DFBA RR53
  • Cleaning – W43/44
  • Collimator installation W44
  • CMS-ZDC optical fiber installation; W44


General planning+Activities:

  • W40: Pressure test of the arc / DSL still disconnected
  • W41-W42:
    • Consolidation
      • All lines under nominal pressure up to Wed. 10/10 - AT-VAC/ACR
      • Cryo-valves reconfiguration Thu.-Fri. 11-12/10 - AT-ACR
      • Opening and repairs of leaks (two leaks found: 7R4 ~2*10^-6 [mbar*l/s] and 7L4 ~10^-5 [mbar*l/s]) - AT-MCS/VAC
    • Checks the jacks - stable pressure conditions needed - 2 days needed - date to be decided, possibly on Friday 12/10 - TS-IC
    • ElQA.DOC from Wed. 10/10 - AT-MEL
    • Manometer installation on the ceiling at the level of RE38 – trolley passage difficult – TS-IC
  • W43: DFBLC-DSLC-DFBAF interconnections start (possibility to start during W42 if compatibility with consolidation)
  • W50: Flushing

Edit, attach file or add comment

Created by Main.AntonioVergara on 2007-10-09 - 11:51

From Pressure Tests to Cooldown - Point 8 - Monday 09.10.2007

Sector 8-1

  • Inner Triplets
    • Right 8: the leak has been located in line C', precisely on a weld in the middle of the flexible. It is currently being replaced by the US collaborators. Once it is corrected and the works in the SSS completed, the inner triplet will be pressurised again for testing. It should be sufficient for this test to have the C' line in the SSS welded, however it is not sure that only one line can be pressurised without affecting the other lines, hence most likely the test will be carried out at the end of the SSS works (wk 47).
    • Left 1: inner triplet closed and being leak-tested.
  • SSS Q17R8
    • The replacement of the SSS started this week and will last till wk 46.
    • Local electrical tests will progress in parallel with the assembly of the SSS.
  • MB1158
    • The diode has been cleaned up. A continuity test will be done before closure beginning of next week. The final validation cannot be done before the TP4-A before the flushing.
  • DFBAs repairs
    • A latest date for taking a decission about repairing or not the DFBAs has to be set for not jeopardizing the starting date of the flushing.
    • Paolo reminds that the time used for reparing the DFBAs will be different for each sector due to the different configurations.
  • ElQA
    • TP4-A is planned for week 47, just before the flushing.
    • Low voltage continuity tests of the dipole lines will be carried ut during the SSS replacement between the DFBAA and C18R8
    • The non-conformities being corrected by INEO this week will be re-checked once they are closed.
  • AUG Test
    • The yearly AUG test for the whole Point 8 is planned for October 19. HCC will study how can we get the maximum profit from it.
  • A4L1
    • Vacuum mechanical works will start on week 43 and will last till week 47.
    • Collimator to be installed week 44.
    • Neg-activation weeks 48 and 49 (overlaps with one week of flushing)
  • LHCb Compensator Magnet Tests
    • The three compensator magnets will be on today and tomorrow. No interference with other activities.
  • Cryogenics
    • Nothing pressurized in the cold circuits.
    • Warm helium line a few mbars above atm pressure
    • Installation of the instrumentation rods in the DFBs of Right 8 will be done tomorrow Wednesday


Edit, attach file or add comment

Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-10-03 - JacekSzkutnik
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login